Calculate the work done when 50.0 g of tin dissolves in excess acid at 1.00 atm and 25°C:

Sn(s)+2H^+ (aq)→ 〖Sn〗^(2+) (aq)+ H_2 (g)

Assume ideal gas behavior.
1
Expert's answer
2020-09-14T08:21:52-0400

In chemistry work is found from a formula:

A=p∆V.

From the law of ideal gas :

pV=nRT, work may be expressed as

A=nRT.

n=m/M.

n=50g/118.71g•mol-1=0.421mol.

R is a gas constant.

Temperature should be expressed in Kelvin's degrees:

T=25+273=298°K.

A=0.421×8.314×298=1043(J).

Answer: 1043 Joules.